OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (anonymous):

m = 1/2 Equation of line with one point and slope is given by: \[\huge \color{green}{y - y_1 = m(x-x_1)}\] So, \(y + 4 = \frac{1}{2}(x - 1)\)